OpenStudy (anonymous):

What is the solution of the equation? -10 + sqrt x+8 = -4

OpenStudy (anonymous):

first, subtract 8 and add 10 from both sides

OpenStudy (anonymous):

that will get you sqrt x = -2. then you square both sides

OpenStudy (anonymous):

and that gets you x = 4.

OpenStudy (anonymous):

but now you have to check it

OpenStudy (anonymous):

and when you do, you will see that 4 is wrong.

OpenStudy (anonymous):

these are the choices, x=4 isnt one A. 28 B. -2 C. 36 D. 44

OpenStudy (anonymous):

because 4 is wrong that means there is no solution. are you sure you typed the problem correctly?

OpenStudy (anonymous):

or is it sqrt(x+8)

OpenStudy (anonymous):

yeah\[-10+\sqrt{x+8}=-4\]

OpenStudy (anonymous):

i thought you meant sqrt of only x not x+8. \[-10 +\sqrt{x+8}=4\] \[-10 +x+8 = 16\] \[x = 16+10-8\] \[x = 28\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem